A 16 gm. mass is moving at 30cm/sec while a 4gm mass is moving at an opposite

direction at 50 cm/sec. They collide head on and stick together. What is their

velocity after collision?


1
Expert's answer
2021-03-14T19:18:54-0400

Let's apply the law of conservation of momentum:


"m_1v_1-m_2v_2=(m_1+m_2)v_f,""v_f=\dfrac{m_1v_1-m_2v_2}{(m_1+m_2)},""v_f=\dfrac{0.016\ kg\cdot0.3\ \dfrac{m}{s}-0.004\ kg\cdot 0.5\ \dfrac{m}{s}}{(0.016\ kg+0.004\ kg)}=0.14\ \dfrac{m}{s}."